OpenStudy (buttermequeasy):

If r(x) = −4x − 1 and t(x) = 2x − 1, find (r + t)(7). −16 16 14 −14

OpenStudy (anonymous):

\[r(7)=-4\times 7-1=-28-1=-29\] \[t(7)=2\times 7-1=14-1=13\]

OpenStudy (anonymous):

and \[-29+13=-16\]

OpenStudy (mertsj):

(r+t) = -2x -2 (r+t)(7) = -2(7)-2 = -16

OpenStudy (anonymous):

-16

OpenStudy (buttermequeasy):

Thanks. c: